Suppose that the annual income of a fast-food chain is normally distributed with a mean of $72,000 and a standard deviation of $9,000. a. What is the probability that a randomly selected fast-food restaurant will generate an annual income between $65,000 and $75,000? (2 Marks) b. What is the probability that a randomly selected fast-food restaurant will generate an annual profit of more than $80,000? (2 Marks) c. What minimum income does a fast-fast restaurant need to earn to be in the top 5% of incomes? (2 Marks) 3 d. What maximum income does a fast-food restaurant need to earn to be in the bottom 30% of incomes? (2 Marks)


1
Expert's answer
2021-04-25T16:07:54-0400

Let x denote the random variable for annual income.

X~N(72000,90002)

a)P(65000<X<75000)

"z_1=\frac{65000-72000}{9000}" =-0.78

"z_2=\frac{75000-72000}{9000}" =0.33

P(-0.78<z<0.33) from the z tables

=0.4116

b)P(X>80000)

"z=\frac{80000-72000}{9000}" =0.89

p(z>0.89) from the z table

=0.1867

c) To be in the top 5% the ; we need the value; P(X<x)=0.95

the corresponding z value is; "\phi ^{-1}(0.95)" =1.6449

"1.6449=\frac{X-72000}{9000}"

X=$86804.1

d) P(X<x)=0.30

"\phi^{-1}(0.3)=-0.5244"

"-0.5244=\frac{X-72000}{9000}"

X=$67280.4